OpenStudy (anonymous):

How do we solve -3(x)+6<-9 and graph it?

OpenStudy (stefrheart):

First you need to simplify it

OpenStudy (anonymous):

So I got x<5

OpenStudy (anonymous):

But like how does this happen : x<5 or -5<x ?

OpenStudy (anonymous):

so you reduce it to -3x < -15, then divide by -3. When you divide or multiply by a negative number, the inequality sign flips. For example -1*3 = -3 < 5, but if you divide both sides by -1 you get 3 on the left and -5 on the right, and so the inequality flips. Now to graph it, you include in the graph all possible values of x. Which in this case, means all numbers that fit > 5.

OpenStudy (anonymous):

-3x<-15 it would actually be x>5 because a negative divided by a negative becomes a positive and when you divide by a negative you are supposed to flip the inverse sign

OpenStudy (anonymous):

And what bout this one 5(x-3)-8>2

OpenStudy (anonymous):

how far are you into it?

OpenStudy (anonymous):

Well I got x>5

OpenStudy (anonymous):

so 5(x-3) > 10, x - 3 > 2, x > 5. that seems solid

OpenStudy (anonymous):

But my teacher said the right answer is 1<x<5 and I don't know why

OpenStudy (anonymous):

so if you plug in a 2 there (which is in the range your teacher describes), you get 5(2-3) - 8 = 5(-1) -8 = -5 -8 = -13 which is not greater than 2.

OpenStudy (anonymous):

Are you sure you're being asked to solve for x?

OpenStudy (anonymous):

or that all the signs are right?

OpenStudy (anonymous):

It just says that we should graph and answer to make the equation true

OpenStudy (anonymous):

What does graphing involve for you guys?